OpenStudy (anonymous):

find f-g given f(x)=6x^2-3 and g(x)=2x^2-5

jimthompson5910 (jim_thompson5910):

f - g = f(x) - g(x) f - g = (6x^2 - 3) - (2x^2 - 5) f - g = 6x^2 - 3 - 2x^2 + 5 f - g = ???

OpenStudy (anonymous):

thanks :)

jimthompson5910 (jim_thompson5910):

you're welcome, tell me what you got

OpenStudy (anonymous):

is it 4x^2+2 ?

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

yay! thanks soo much for helping :)